OpenStudy (anonymous):

Solve the problem by stating the equation you would use to solve and then solving it. Be sure to show all of your work. After spending $3850 for tables and $1750 for chairs, a convention center manager finds that 30% of his original budget remains. Find the amount that remains. Round to the nearest dollar.

OpenStudy (anonymous):

i think its this: 3850 + 1750 • 30% = x but I am not sure.

OpenStudy (dumbcow):

If 30% remains, then he spent 70% on tables and chairs 0.7x = 3850 + 1750

OpenStudy (anonymous):

so is 8,000 my answer?

OpenStudy (anonymous):

lol I still don't know what im doing

OpenStudy (dumbcow):

sorry, no 8000 would be the budget, its asking for what the 30% remaining is so just find 30% of 8000

OpenStudy (anonymous):

oh ok so the original set up you gave me will give me the budget amount which is 8,000 and then my final answer is 30% of that?

OpenStudy (dumbcow):

exactly

OpenStudy (anonymous):

I guess my issue is that I need to set up the equation to just give me the 30% I don't know think my instructor will accept 2 equations. Thanks for your response!

OpenStudy (dumbcow):

why does it matter as long as you solve the problem? you could just substitute one equation into the other x = (3850+1750)/0.7 30% budget = (0.3/0.7)*(3850+1750)

OpenStudy (anonymous):

that works thank you

OpenStudy (anonymous):

\[0.3\left(\frac{3850 + 1750}{1-0.3}\right)=\text{$\$$2400} \]

OpenStudy (anonymous):

oh wow thanks this set up look better to me but how did you get 2400?

OpenStudy (anonymous):

Original budget amount was:\[\left(\frac{3850 + 1750}{1-0.3}\right)=\left(\frac{3850 + 1750}{0.7}\right)=8000 \]\[30\% \text{ of } 8000=\frac{30}{100}8000=2400\]

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

thank you for the help robtobey and you too dumbcow